A Bibliography of the Biennial IEEE Hot Topics in Operating Systems Workshops

Nelson H. F. Beebe University of Utah

Department of Mathematics, 110 LCB

155 S 1400 E RM 233 Salt Lake City, UT 84112-0090

USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

E-mail: beebe@math.utah.edu, beebe@acm.org,
beebe@computer.org (Internet)
WWW URL: http://www.math.utah.edu/~beebe/

30 December 2016 Version 1.04

Title word cross-reference

-Kernel [LIJ97].

3 [OS89].

64-bit [CLBHL92].

Abstraction [CRL01, MST93, CBG^{+}92]. Abstractions [EK95, Puc92]. Access [JELP99a, JELP99b, KL99]. Accountable [RPEM99]. Accounting [NM01]. Achieved [LESH97]. Achieving [CJT01]. across [Bar93a]. Active [LPTL99, RL01, LMMW89, BEW^{+}01]. Adaptation [YRP99]. Adapting [SS97]. Adaption [BCHS01, dLWZ01]. Adaptive [BEW^{+}01, Jon93, LYI95, OS95, SM99]. Address [CFL93, Che93, CVP99, GKB01, LRD95, CCJZ92, CLBHL92]. after [WRPK01]. Agent [vNBTvS01]. Agents [JVS95, LJJM01]. Agile [RL01]. AIX [OS89]. algorithms [Atk89]. Amoeba [Mul87]. amount [Mog89]. Answering [Ham95]. Anypoint [YCV01]. Appliance [KA01]. Application [BM01, EGK95, KS93, Orr95, SKSZ99, Wat95, And92]. Application-Level [EGK95]. Application-Specific [BM01, KS93, SKSZ99, And92]. Applications [MK97, Sta95, YLS^{+}01, CB89, Jef92, LMMW89]. Applying [LW97, POBF01]. Approach [BEW^{+}01, GRC97, JF95, MMWV01, PF93, RBP01, SM99, Wil92]. April [IEE92].

Architecture [BBM89, RBS99, SMBNI01, CLBHL92]. Architectures [dLWZ01]. Area [WA93]. Arias [PCDH95]. Arizona [IEE99]. Arm [EH01]. Aspect [CKF^{+}01]. Aspect-Oriented [CKF^{+}01]. Assistants [LMR93]. Astray [DPZ97, Mog92]. at-most-once [LSW89]. Attacks [LIJ97]. Authentication [Gan01]. Authored [Jon93]. autonomy [Red92]. Availability [HBM^{+}89]. Available [SSI01, Sat89]. AVM [EGK95]. Awareness [CEM01].

Backup [KA01]. Bandwidth [ZD95]. Based [LMR93, PF93, Raw97, VHT95, BDD^{+}89, CCJZ92, LSW89, PC99]. batch [LL92]. Be [CKK95, Day93, JR99, DPH92, SG92]. Benchmarking [SKSZ99]. benefits [HF92]. Better [CN01, Day93, GN01, NLP99]. Between [How93]. Beyond [GKB01]. Big [Mog93]. Biscayne [IEE92]. bit [CLBHL92]. Bossa [BM01]. bottleneck [LS92]. Boundaries [Bar93a]. Box [BA99]. Brittle [Mog99]. Buffer [AYC99]. Building [CFH^{+}93, DBK^{+}01, FSA97, GRC97, SN97].

CA [IEE89, IEE93]. Caches [GRC97, POBF01]. Caching [GS95, LPTL99, NK93, Hau92]. Cambridge [IEE87]. Capacity [KA01, MST93]. Cape [IEE97]. Careful [MP93]. Case [ABKM01, AYC99, BBIM93, Dra93, Ell99a, Ell99b, GS95, KL99, NLP99, SCA99, SKSZ99, And92, Wil92]. CE [TI99]. Censor [RL01]. Censor-Resistant [RL01]. Censorship [WM01]. Central [Fra97]. Chaining [Bar93a]. Choices [CT95]. Chord [DBK^{+}01]. Client [Ing99]. Client-Server [Ing99]. Clients [BBIM93, MSK^{+}93]. clocks [LSW89]. Cluster [PC99]. Cluster-based [PC99]. Clusters [CD01]. Cod [IEE97]. Coda [SKS87, Sat89]. Code [Fra97, KMF01, Jon92]. Command [TRP99a, TRP99b]. Comments [Hag89]. Commercial [Ham95]. commonplace [SG92]. Communication [KBK99, LHR01, Wat95, YCV01, ZD95, And89]. Communications [KGLR92, MMOP95, RH89]. Communications-intensive [KGLR92]. Communications-Oriented [MMOP95]. Compiler [HM99]. Complex [DMK95, Gri01]. Component [Men97]. Components [FLCV97, HAW^{+}01]. Composed [KA01]. Composition [Jon93, KMF01]. computation [BDD^{+}89]. Computer [FSA97]. Computers [CDLM93, GT93, How93, Wil92]. Computing [BCK93, CEM01, FNS01, GDH^{+}01, HJM^{+}01, LK93, LS95, MH87, RBP01, SN97, SAG^{+}93, vRTS87, DM92, Duc92]. Conditioned [WC01]. Conductor [YRP99]. Confidences [Gan01]. Conflict [Mye93]. Congestion [RBS99]. Connected [How93]. connector [LMMW89]. Conquest [WRPK01]. considerations [RH89]. Considered [WC01]. Consistency [LS95, PCDH95, HBM^{+}89]. Consistent [JF95]. constructing [LMMW89]. Continuous [Sta95]. Control [JELP99a, JELP99b, Sta95, Puc92]. cooperation [CBG^{+}92]. Coordinated [BCHS01]. Cost [EH01]. Critical [Jon93]. Cryptography [Bla93]. Cryptography-in-the-Large [Bla93]. Currencies [SHS99]. Customizable [KLM^{+}93]. Customization [AFG^{+}97].

Data [BOKT99, LS95, Mog01, OS95, VRTW01]. Data-Intensive [BOKT99]. Dawn [WGJ^{+}01]. dead [Red92]. decoupled [DPH92]. Denial [LIJ97]. Denial-of-Service [LIJ97]. Design [LYI95, PGC01, WC01, HPR89, Pas92].

Desktop [Mog93]. Developing [MMWV01]. Development [Raw97]. Device [Puc92]. Devices [GN01, LMMW89]. Digital [LMR93]. DiPS [MMWV01]. Directions [DBFF97, GDH^{+}01, WC01]. Disconnected [NK93, PMDD01]. Disk [GPS92, KA01, LS92]. Disks [WRPK01]. Distributed [BEW^{+}01, CCJZ92, CKK95, Chi97, DBK^{+}01, GC93, JF95, MH87, vNBTvS01, SKS87, WH01, YRP99, Atk89, BDD^{+}89, BA92, CBG^{+}92, HBM^{+}89, NY92, Sat89, SKKM92, WBD^{+}89]. Divergence [RLB01]. Diverse [FSA97]. Document [WM01]. Documents [LPTL99]. domain [Wel92]. Don’t [MS01]. DOORS [PMDD01]. DRAM [GPS92]. Drawing [BH99]. Drives [KA01]. DSL [BM01]. During [SS95]. Dynamic [SS95].

easy [Kaz89]. Echo [HBM^{+}89]. Economic [SM99]. Economy [WGJ^{+}01]. Effective [Wat95]. Efficient [ABSK95, Atk89, CFH^{+}93, KL99, LSW89]. Eighth [IEE01]. Elephant [SFHV99]. Elmau [IEE01]. Embedded [SMBNI01]. Enabled [MBL01]. End [RBS99]. End-System [RBS99]. Energy [CD01, NM01]. Entanglements [WM01]. environment [GM89, LL92, Sat89, Sch89]. Environments [Men97, ZD95, DM92]. eNVy [WZ93]. eOS [WGJ^{+}01]. Equipment [TI99]. Escaping [LS92]. Estimation [NLP99]. evaluate [BDF92]. Event [CJT01]. Execution [FNS01, RPEM99]. Experience [BDD^{+}89, Raw97]. Extensibility [EH01]. Extensible [DPZ97, GB97, VHT95]. Extensions [SHS99]. Exterminate [EK95].

Fail [ADAD01]. Fail-Stutter [ADAD01]. Failure [Fle97, WBD^{+}89]. fast [LS92]. Fault [ADAD01, CRL01, BBM89]. fault-tolerant [BBM89]. Feedback [PF93]. Feedback-Based [PF93]. Fifth [IEE95].

File [KL99, KS93, MS01, SFHV99, SKS87, TI99, WA93, WRPK01, BA92, HBM^{+}89, Mog89, Sat89, SKKM92]. Files [BA99, How93]. Filtering [ZD95]. FLEX [CFH^{+}93]. Flexibility [GKB01]. Flexible [CFH^{+}93, JELP99a, JELP99b, MK97, POBF01]. Florida [IEE92]. Flux [FLCV97]. Forgets [SFHV99]. Formal [TTML97]. foundation [RBF^{+}89]. Fourth [IEE93]. Framework [BM01, YRP99]. Fresh [Cha97]. friendly [LL92]. Futz [HJM^{+}01].

Games [Chi97]. General

[Ing99, PKL99, ZD97]. General-Purpose [PKL99]. Generalize [Fle97]. Generation [Fra97, TRP99a, TRP99b]. Generic [NK93, PCDH95]. Geographical [GS95]. Germany [IEE01]. Global [CJT01, VRTW01, And89, CCJZ92]. Good [WH01]. Goodnes [PNG^{+}01]. Got [HM99]. Grand [LRD95]. Groups [Day93]. Groupware [Day97]. Grove [IEE89]. GUIDE [BDD^{+}89].

Handling [BBIM93]. Hardware

[BF97, CVP99]. Harmful [WC01]. Harvest [FB99a, FB99b]. Having [JR99]. Headed [CKK95]. Heavyweight [Bar93a]. Herald [CJT01]. HeRMES [MBL01]. Hey [HM99]. Hierarchies [CL99a, CL99b]. Hierarchy [Cha97]. High [KA01, MBL01, SS95]. High-Capacity [KA01]. High-Performance [MBL01]. Higher [Ell99a, Ell99b]. Higher-Level [Ell99a, Ell99b]. Highly [MP93, SSI01, Sat89]. Highly-Available [SSI01]. Hinting [PNG^{+}01]. Host [Joh93]. Hot [HAW^{+}01, IEE95, IEE97, IEE99, IEE01]. Hot-Swappable [HAW^{+}01]. HotOS [IEE95, IEE99, IEE01]. HotOS-V [IEE95]. HotOS-VII [IEE99]. HotOS-VIII [IEE01]. HTTP [Mog01].

I/O [Pas92, SS95]. Ideas [LUEJ99]. IEEE [Cab90]. II [IEE89, Ano89]. Implementation [FLCV97]. Implementors [TTML97]. Implications [CDLM93]. Importance [WH01]. Improve [CRL01, Orr95]. Improving [LS95]. Independently [Jon93]. Indirect [BBIM93]. information [Col92, Sch89]. Informed [SCA99]. Infrastructure [WH01, YLS^{+}01]. Input [TRP99a, TRP99b]. integrating [DM92]. Integrity [Mye93]. Integrity/Performance [Mye93]. Intelligent [ZD95]. Intensive [BOKT99, KGLR92, Pas92]. Interact [PC99]. Interaction [BBIM93]. Interface [LK93, LHR01, And89, Jon92, RH89]. interfaces [Hau92]. Internet [SBK01, GRC97, LW97]. Internetworking [Joh93]. interposing [Jon92]. Interprocess [LHR01]. Intra [CVP99]. Intra-Address [CVP99]. Introspective [BOKT99]. Invocation [Bar93a]. IPC [HKW93, JELP99a, JELP99b, LESH97]. Island [IEE95]. Isolation [LS95]. Isolation-Only [LS95]. Issue [BSPB95, Red92]. Issues [CFL93, Duc92, HJM^{+}01, LK93, PGC01, TDW93]. ISTORE [BOKT99]. IV [IEE93].

Java [BH99]. jobs [LL92]. Just [NM01, vRTS87]. JX [GKB01].

Kernel [Dra93, LIJ97, POBF01, VHT95, HPR89, Jef92]. Kernels [DPZ97]. Key [IEE92].

labyrinth [OM92]. Lana [RBP01]. language [SLM89]. Large [Bla93, DR01, GRC97, SMBNI01, CB89]. Large-Scale [DR01]. Latencies [SS95]. Latency [JR99, GPS92]. Layer [SSI01]. Lazy [LW01]. Leading [DPZ97, Mog92]. Leg [EH01]. Level [Ell99a, Ell99b, EGK95]. library [Wel92]. Life [WRPK01]. Lightweight [Bar93a]. like [FS89]. Limited [ZD95]. Line [BH99]. Links [ABSK95]. Lite [AFG^{+}97]. LITTLE [HHRB92]. logging [SKKM92]. Longer [Nil95]. Look [Cha97]. Lookup [DBK^{+}01]. Lost [OM92]. Lottery [SHS99].

Mach [RBF^{+}89]. made [Kaz89]. Main [WZ93]. Making [LL92]. Management [Cha97, CD01, Ell99a, Ell99b, GKB01, Jon93, JLDB95, KBK99, OS95, RBS99, SM99, TRP99a, TRP99b, VRTW01]. Managing [MST93]. Manipulating [Bar93a]. Mansion [vNBTvS01]. March [IEE99]. Mass [WA93]. Massachusetts [IEE87, IEE97]. massive [BA92]. May [Day93, IEE95, IEE97, IEE01, JR99]. Mechanism [DMK95]. mechanisms [LMMW89]. Medium [Hag92]. Memories [Mog93]. Memory [CKK95, Cha97, EGK95, GC93, JF95, KTNW93, LUEJ99, WZ93, Atk89, CBG^{+}92, Hag89, Hag92, NY92, SG92]. mesh [Col92]. Messages [SMBNI01, LSW89]. Meta [DMK95, Orr95]. Meta-Protocols [Orr95]. Meta-Scripts [DMK95]. Methods [Che93, OS95, TTML97]. Metrics [Mog99]. microbenchmarks [BDF92]. Microkernel [LMR93, Raw97]. Microkernel-Based [LMR93, Raw97]. Middleware [CEM01]. Millennium [DBFF97]. Mistrust [Bla93]. Mobile [BBIM93, BCK93, CDLM93, CEM01, JVS95, Joh93, LK93, LJJM01, LS95, MSK^{+}93, SN97, SAG^{+}93, Duc92]. Mobility [LJJM01, SBK01, DM92]. Model [Mog01, Neu89]. Modeling [KL99]. Modelling [RLB01]. Modular [JLDB95]. Modularity [DPH92]. Modulation [SN97]. Modules [Dra93]. Monitoring [SS97]. most [LSW89]. MRAM [MBL01]. MRAM-Enabled [MBL01]. Multi [vNBTvS01, Raw97, SHS99, SLM89]. Multi-Agent [vNBTvS01]. multi-language [SLM89]. Multi-Resouce [SHS99]. Multi-Server [Raw97]. multi-user [SLM89]. Multimedia [CT95, KKS95, Sta95, Jef92, LMMW89, Pas92]. Multiple [KTNW93]. multiprocessing [FS89]. Multiprocessor [MH87, BBM89]. My [HM99].

Napa [IEE93]. Need [KLM^{+}93, MK97]. Needn’t [EH01]. Needs [Day97, OS95]. Nemesis [NM01]. Network [BF97, BOKT99, MSK^{+}93, SSI01, RH89]. Networked [BCHS01, Chi97, GT93]. Networks [ABSK95, ABKM01, KBK99, MP93, RL01, SMBNI01, CB89]. Neutrality [LUEJ99]. Never [SFHV99]. Next [DBFF97, TRP99a, TRP99b]. Next-Generation [TRP99a, TRP99b]. Niche [Nil95]. No [HJM^{+}01, Nil95]. No-Futz [HJM^{+}01]. NonVolatile [WZ93]. Notification [CJT01, Day97]. November [IEE87]. NT [JR99].

O [Pas92, SS95]. Object [CT95, Day93, GC93, VHT95, BDD^{+}89, GM89]. Object-Based [VHT95, BDD^{+}89]. Object-Oriented [CT95, GM89]. Objects [Bar93a]. Occasionally [How93]. October [IEE93]. once [LSW89]. Only [LS95]. Opal [CLBHL92]. open [HPR89, RBF^{+}89, SLM89]. Operating [BF97, Cab90, CDLM93, CT95, DBFF97, EK95, GKB01, GM89, Ham95, HM99, IEE87, IEE92, IEE93, IEE95, IEE97, IEE99, IEE01, JVS95, JLDB95, KKS95, KLM^{+}93, LYI95, LMR93, MK97, Men97, Mog99, MMOP95, Mul87, PC99, Raw97, SS97, TDW93, And92, BDD^{+}89, CCJZ92, FS89, Hag89, HPR89, IEE89, OS89, Puc92, RBF^{+}89, SLM89, Wil92]. Operation [NK93, PMDD01]. Optimistically [KS93]. Orcas [IEE95]. Oriented [CT95, CKF^{+}01, MMOP95, GM89]. ought [SG92]. Overcome [SS95]. Overlay [ABKM01].

Pacific [IEE89]. Page [KTNW93]. Pages [Day93]. parallel [Atk89]. Paramecium [VHT95]. ParcTab [SAG^{+}93]. Partitioned [YLS^{+}01]. Partitioning [Wat95]. PAST [DR01]. Pattern [KL99]. PDAs [TDW93]. Peer [DBK^{+}01, DR01]. Peer-to-Peer [DBK^{+}01, DR01]. Penalty [HKW93]. Performance [GKB01, HKW93, LESH97, MBL01, Mye93, BDF92, RH89]. Persistent [DR01, HKW93]. Personae [BCK93]. Personal [LMR93, MH87, Wil92]. Personalities [Fle97]. Pervasive [FNS01, GDH^{+}01]. Pinning [LUEJ99]. Placement [VRTW01]. plea [Hau92]. Plumbing [WH01]. Plumbing-Reconsidering [WH01]. PLURIX [FS89]. Policies [BM01]. Portable [GT93]. Position [BM87, WBD^{+}89]. Power [Ell99a, Ell99b]. Practical [IH99a, IH99b, TTML97]. presentation [Ano89]. Preventing [LIJ97]. Pricing [NM01]. Probabilistic [RLB01]. Problems [JR99, ZD97]. proceedings [IEE89, IEE95, IEE87, IEE92]. Process [LW01]. Processes [LUEJ99]. Processing [KKS95, LS92]. Processor [MST93]. Program [GT93]. Programmable [PC99, PGC01]. Programs [RPEM99]. project [HHRB92, NC89]. Properties [LPTL99]. proposal [Wel92]. PROST [Ano89]. protected [CCJZ92]. Protection [BSPB95, CVP99, GKB01, DPH92]. Protium [YLS^{+}01]. Protocol [YCV01]. Protocols [MP93, Orr95, RL01, SCA99]. prototype [Sch89]. Providing [MSK^{+}93]. Provisional [LUEJ99]. Proxies [ZD97]. public [Wel92]. Publishing [WM01]. Purpose [Ing99, PKL99, ZD97]. Push [GS95]. Push-Caching [GS95].

Query [LW97]. Questions [Ham95].

Raven [NC89]. reads [GPS92]. Real [CN01, Ing99, Jon93, JLDB95, KKS95, LHR01, Nil95, Sta95, Jef92]. Real-Time [Jon93, JLDB95, KKS95, LHR01, Nil95, Sta95, Jef92]. rebalancing [HF92]. Reboot [CF01]. Reconciliation [How93]. Reconsidering [SBK01, WH01]. recovery [WBD^{+}89]. Recursive [CF01]. Recycle [GRC97]. Red [BH99]. Redirection [JELP99a, JELP99b]. Reduce [GRC97]. Reducing [HKW93]. Reflection [LYI95]. Relevance [HKW93]. Reliability [CL99a, CL99b]. Reliable [JF95, MBL01, MP93]. Remote [FNS01, NY92]. Replaceable [Dra93]. Replacement [BA99, Hag92]. Replica [RLB01]. Replicated [KS93, Kaz89]. Research [DPZ97, Ham95, HJM^{+}01, Mog99]. Reserves [MST93]. Resilient [ABKM01, SKS87]. Resistant [RL01, WM01]. Resolution [KS93]. Resolving [Mye93]. Resouce [SHS99]. Resource [GKB01, Jon93, JLDB95, NM01, SM99, WGJ^{+}01, vRTS87, NY92]. Restartability [CF01]. Results [JR99]. Reusable [FLCV97]. Reuse [GRC97]. Revisited [SHS99]. Rialto [JLDB95]. Rico [IEE99]. Right [Ham95, Mog89]. Rigorous [Mog01]. Rio [IEE99]. Robustness [Gri01]. Role [SN97]. Routers [PKL99, PC99, PGC01]. Routing [LW97]. Run [Dra93, Fra97]. Run-Time [Dra93, Fra97].

Safe [GKB01]. Sake [PNG^{+}01]. Scalable [FB99a, FB99b]. Scale [DR01, CB89]. Scalpel [CF01]. Schedule [LUEJ99]. Scheduling [BM01, Ing99, PF93, SHS99]. Schloss [IEE01]. Scout [MMOP95]. Scripts [DMK95]. Search [LW97, SS95]. Second [IEE89, Cab90]. Secure [EH01, MK97]. Security [GN01, GB97, IH99a, IH99b]. Segmentation [CVP99]. Self [Bar93b, FNS01, SS97]. Self-Adapting [SS97]. Self-Monitoring [SS97]. Self-Tuned [FNS01]. Self-Tuning [Bar93b]. Separating [LJJM01]. September [IEE89]. Server [CD01, Ing99, MS01, Raw97, Mog89]. Servers [AYC99, MB93, Kaz89]. Service [CJT01, DBK^{+}01, Fra97, KMF01, LIJ97, LUEJ99, MB93, MSK^{+}93, HF92]. Service-Neutrality [LUEJ99]. Services [BOKT99, Day97, DMK95, Jon93, Orr95, SSI01, WC01]. Sets [SS95]. Seventh [IEE99]. Share [How93]. Shared [CKK95, GC93, JF95, OS95, SG92, Atk89, CBG^{+}92]. Should [CKK95, DPH92]. Single [CFL93, CLBHL92]. Sixth [IEE97]. Sizes [KTNW93]. Sledgehammer [CF01]. Small [Nil95]. Smart [SMBNI01]. Smartcards [IH99a, IH99b]. Smarter [GN01]. Soft [Ing99]. Software [Bar93b, BSPB95, Che93, HAW^{+}01, Men97, MMWV01]. Solid [CDLM93]. Solid-State [CDLM93]. Solved [ZD97]. Some [CFL93]. Space [CFL93, CVP99, CCJZ92, CLBHL92]. Spaces [GKB01, LRD95]. Spaces-Flexibility [GKB01]. Specialized [Nil95]. Specific [BM01, KS93, SKSZ99, And92]. SPECmarks [Mog92]. Sprite [WBD^{+}89]. State [CDLM93, WBD^{+}89]. statelessness [Mog89]. statement [BM87, WBD^{+}89]. stations [Wil92]. Storage [BOKT99, DR01, MBL01, VRTW01, WA93, WZ93]. Streaming [LHR01]. Streams [BEW^{+}01]. Structure [CKF^{+}01]. Study [JR99]. Stutter [ADAD01]. Summary [Cab90]. Support [Bla93, Chi97, JVS95, KTNW93, NK93, PCDH95, PKL99, PC99, PGC01, Sta95, SSI01, GM89, Hau92, Jef92]. Supporting [BCHS01, Col92, CB89, HAW^{+}01, Jon93, KS93, PMDD01]. Swappable [HAW^{+}01]. Switching [LW01]. Synchronization [PCDH95]. synchronized [LSW89]. Synchronous [Day97]. System [BF97, CDLM93, CT95, Che93, CKF^{+}01, DBFF97, EK95, Fra97, GKB01, HM99, HAW^{+}01, JVS95, JLDB95, KS93, LYI95, LMR93, MH87, MMWV01, MMOP95, Mul87, vNBTvS01, PGC01, RBS99, Raw97, SFHV99, SKS87, SAG^{+}93, SMBNI01, TDW93, TRP99a, TRP99b, WA93, WRPK01, WZ93, BDD^{+}89, BDF92, CLBHL92, FS89, GM89, HBM^{+}89, HPR89, Jon92, Neu89, OS89, Pas92, Puc92, Sat89, SLM89]. Systems [Bar93b, BCHS01, BEW^{+}01, CFH^{+}93, CFL93, Chi97, DBK^{+}01, FSA97, FB99a, FB99b, Gri01, GB97, GDH^{+}01, Ham95, IEE87, IEE92, IEE93, IEE95, IEE97, IEE99, IEE01, Ing99, IH99a, IH99b, KKS95, KLM^{+}93, dLWZ01, LW97, MK97, Men97, Mog99, PC99, SN97, SS97, SMBNI01, TI99, WH01, And92, Atk89, BA92, CCJZ92, Hag89, IEE89, NY92, RBF^{+}89, SKKM92, Wel92, Wil92, Cab90].

TCP [ABSK95]. Technical [Cab90]. Techniques [HKW93]. technology [OS89]. term [Hag92]. Terminal [vRTS87]. testing [Atk89]. Theory [LRD95]. Think [JR99]. Thinking [LW97]. Third [IEE92]. Throughput [NLP99]. Tickets [SHS99]. Time [Dra93, Fra97, Ing99, Jon93, JLDB95, KKS95, LHR01, Nil95, Sta95, Jef92]. Time-Critical [Jon93]. Tolerance [ADAD01, CRL01]. Tolerant [FB99a, FB99b, BBM89]. Tool [CFH^{+}93, TTML97]. Toolbox [PF93]. Toolkit [FLCV97]. Topaz [MH87]. Topics [IEE95, IEE97, IEE99, IEE01]. Trace [SN97]. Tracing [Che93]. tradeoffs [HBM^{+}89]. transaction [LS92]. Transactions [LS95]. Transparency [CEM01]. Transparent [Bla93, KBK99]. Transparently [Jon92]. Transport [SCA99, SSI01]. Trust [MS01]. Tuned [FNS01]. Tuning [Bar93b]. Turning [CF01]. Type [GKB01, Mog01]. Type-Safe [GKB01]. typed [CBG^{+}92].

Ubik [Kaz89]. Ubiquitous [Joh93].

ubiquity [SKKM92]. Unified

[LRD95, RBS99]. Unifying [MMWV01]. Unix [FS89]. Unix-like [FS89]. Unlimited [LUEJ99]. Unsolved [ZD97]. Untrusted [LUEJ99, RPEM99]. Usage [MST93]. Use [MP93, WBD^{+}89]. User [LK93, TRP99a, TRP99b, Jon92, SLM89]. Using [BDF92, CRL01, How93, LYI95, LS95, SS95, CVP99, JELP99a, JELP99b]. Utility [DR01].

V [IEE95]. via [GN01]. Video [MSK^{+}93]. VII [IEE99]. VIII [IEE01]. Virtual [CN01, EGK95, KTNW93, CCJZ92, Hag89, Hag92, Neu89]. Virtualization [WC01]. vs [CEM01]. VVM [POBF01].

Washington [IEE95]. waystation [DM92]. Web [DMK95, POBF01]. WebOSes [LIJ97]. Well [WC01]. Well-Conditioned [WC01]. Where [CKK95]. Wide [WA93]. Windows [JR99]. Wireless [ABSK95, KBK99, RBP01, Wat95, Duc92]. without [MB93]. work [Wil92, HHRB92]. Workshop [Cab90, IEE87, IEE89, IEE92, IEE93, IEE95, IEE97, IEE99, IEE01]. Workstation [Cab90, CBG^{+}92, IEE87, IEE89, IEE92, IEE93, Mul87, Red92, vRTS87, And89, DM92, Hag89, Puc92, RH89, Sat89, IEE89]. Workstations [Neu89, BBM89, CB89, KGLR92, LL92, OM92]. WWOS [IEE89, IEE93, Ano89]. WWOS-II [IEE89, Ano89]. WWOS-IV [IEE93].

x [HPR89]. x-kernel [HPR89].

Xenoservers [RPEM99]. xFS [WA93].

Yield [FB99a, FB99b]. You’re [JR99].

Zero [KMF01]. Zero-Code [KMF01].

References

Andersen:2001:CRO

[ABKM01] David Andersen, Hari Balakrishnan, M. Frans Kaashoek, and Robert Morris. The case for resilient overlay networks. In IEEE [IEE01], pages 152–160. ISBN 0-7695-1040-X. US$135.00. URL http:

//nms.lcs.mit.edu/papers/
ron-hotos2001.pdf. IEEE cat-

alog number PR01040.

Amir:1995:ETN

[ABSK95] E. Amir, H. Balakrishnan, S. Seshan, and R. H. Katz. Efficient TCP over networks with wireless links. In IEEE [IEE95], pages 35–41. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Arpaci-Dusseau:2001:FSF

[ADAD01] R. H. Arpaci-Dusseau and A. C.

Arpaci-Dusseau. Fail-stutter fault tolerance. In IEEE [IEE01], pages 33–40. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Auslander:1997:CL

[AFG^{+}97] Marc A. Auslander, Hubertus Franke, Benjamin Gamsa, Orran Krieger, and Michael Stumm. Customization lite. In IEEE [IEE97], pages 43– 48. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Anderson:1989:WGC

[And89]  David P. Anderson.  The

workstation as global communication interface. In IEEE [IEE89], pages 6–12. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109259.pdf. IEEE

catalog number 89TH0281-6.

Anderson:1992:CAS

[And92] T. E. Anderson. The case for application-specific operating systems. In IEEE [IEE92], pages 92–94. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275682.pdf. IEEE

catalog number 92TH0413-5.

Anonymous:1989:WIP

[Ano89] Anonymous. WWOS-II presentation of PROST. In IEEE [IEE89], pages 101–104. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109276.pdf. IEEE

catalog number 89TH0281-6.

Atkins:1989:ESM

[Atk89]  M. Stella Atkins.  Efficient

shared memory for testing parallel algorithms on distributed systems. In IEEE [IEE89], pages 13–15. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109260.pdf. IEEE

catalog number 89TH0281-6.

Anderson:1999:CBS

[AYC99] D. Anderson, K. Yocum, and J. Chase. A case for buffer servers. In IEEE [IEE99], pages 82–89. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Blaze:1992:TMD

[BA92] M. Blaze and R. Alonso. Toward massive distributed file systems. In IEEE [IEE92], pages 48–51. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275690.pdf. IEEE

catalog number 92TH0413-5.

Ballesteros:1999:BRF

[BA99] F. Ballesteros and S. Arevalo.

The Box: a replacement for files. In IEEE [IEE99], pages 24–31. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Barrera:1993:ICM

[Bar93a] Joseph S. Barrera III. Invocation chaining: Manipulating lightweight objects across heavyweight boundaries. In IEEE [IEE93], pages 191– 193. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348150.pdf. IEEE

catalog number 93TH0553-8.

Barrera:1993:STS

[Bar93b] Joseph S. Barrera III. Selftuning systems software. In IEEE [IEE93], pages 194– 197. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348149.pdf. IEEE

catalog number 93TH0553-8.

Badrinath:1993:HMC

[BBIM93] B. R. Badrinath, Ajav V. Bakre, Tomasz Imielinski, and R. Marantz. Handling mobile clients: a case for indirect interaction. In IEEE [IEE93], pages 91–97. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348166.pdf. IEEE

catalog number 93TH0553-8.

Banatre:1989:AFT

[BBM89] J. P. Banâtre, M. Banâtre, and G. Muller. Architecture of fault-tolerant multiprocessor workstations. In IEEE [IEE89], pages 20–24. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109262.pdf. IEEE

catalog number 89TH0281-6.

Bridges:2001:SCA

[BCHS01] P. G. Bridges, W.-K. Chen, M. A. Hiltunen, and R. D. Schlichting. Supporting coordinated adaption in networked systems. In IEEE [IEE01], page 162. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Banerji:1993:MCP

[BCK93] Arindam Banerji, David L.

Cohn, and Dinesh C. Kulkarni. Mobile computing personae. In IEEE [IEE93], pages 14–20. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348179.pdf. IEEE

catalog number 93TH0553-8.

Balter:1989:EOB

[BDD^{+}89] R. Balter, D. Decouchant, A. Duda, A. Freyssinet, S. Krakowiak, M. Meysembourg, M. Riveill, C. Roisin, X. Rousset de Pina, R. Scioville, and G. Vandome. Experience with objectbased distributed computation in the GUIDE operating system. In IEEE [IEE89], pages 16–19. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109261.pdf. IEEE

catalog number 89TH0281-6.

Bershad:1992:UME

[BDF92] B. K. Bershad, R. P. Draves, and A. Forin. Using microbenchmarks to evaluate system performance. In IEEE [IEE92], pages 148–153. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275671.pdf. IEEE

catalog number 92TH0413-5.

Bustamante:2001:ASA

[BEW^{+}01] F. E. Bustamante, G. Eisenhauer, P. Widener, K. Schwan, and C. Pu. Active Streams — an approach to adaptive distributed systems. In IEEE [IEE01], page 163. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Ballesteros:1997:NHO

[BF97] F. Ballesteros and L. Fernandez. The network hardware is the operating system. In IEEE [IEE97], pages 32–37. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Back:1999:DRL

[BH99] G. Back and W. Hsieh. Drawing the red line in Java. In IEEE [IEE99], pages 116– 121. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Blaze:1993:TMS

[Bla93]  Matt Blaze.  Transparent

mistrust: OS support for cryptography-in-the-large. In IEEE [IEE93], pages 98–102. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL

http://ieeexplore.ieee.org/
iel2/918/8054/00348165.pdf.

IEEE catalog number 93TH0553-8.

Borg:1987:PS

[BM87] Anita Borg and Jeffrey C.

Mogul. Position statement. In IEEE [IEE87], page ?? ISBN ???? LCCN ????

Barreto:2001:BDF

[BM01] L. P. Barreto and G. Muller.

Bossa: a DSL framework for application-specific scheduling policies. In IEEE [IEE01], page 161. ISBN 0-7695-1040-X. US$135.00. URL http://

                                        computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Brown:1999:IIS

[BOKT99] A. Brown, D. Oppenheimer, K. Keeton, and R. Thomas. ISTORE: Introspective storage for data-intensive network services. In IEEE [IEE99], pages 32–37. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Bershad:1995:PSI

[BSPB95] B. N. Bershad, S. Savage, P. Pardyak, and D. Becker. Protection is a software issue. In IEEE [IEE95], pages 62–65. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Cabrera:1990:TSS

[Cab90] Luis-Felipe Cabrera. Technical summary of the Second IEEE Workshop on Workstation Operating Systems. Operating Systems Review, 24(3):7–21, 1990. CODEN OSRED8. ISSN 0163-5980.

Cooper:1989:SLS

[CB89] Robert Cooper and Ken Birman. Supporting large scale applications on networks of workstations. In IEEE [IEE89], pages 25–28. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109263.pdf. IEEE

catalog number 89TH0281-6.

Cohn:1992:WCT

[CBG^{+}92] D. L. Cohn, A. Banerji, P. M.

Greenwalt, M. R. Casey, and D. C. Kulkarni. Workstation cooperation through a typed distributed shared memory abstraction. In IEEE [IEE92], pages 70–74. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275686.pdf. IEEE

catalog number 92TH0413-5.

Carter:1992:DOS

[CCJZ92] J. B. Carter, A. L. Cox, D. B.

Johnson, and W. Zwaenepoel. Distributed operating systems based on a protected global virtual address space. In IEEE [IEE92], pages 75–79. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL

http://ieeexplore.ieee.org/
iel2/891/6831/00275685.pdf.

IEEE catalog number 92TH0413-5.

Chase:2001:EMS

[CD01] J. Chase and R. Doyle. Energy management for server clusters. In IEEE [IEE01], page 165. ISBN 0-7695-1040-X. US$135.00. URL http://

                                        computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Caceres:1993:OSI

[CDLM93] Ramón Cáceres, Fred Douglis, Kai Li, and Brian Marsh. Operating system implications of solid-state mobile computers. In IEEE [IEE93], pages 21–27. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348178.pdf. IEEE

catalog number 93TH0553-8.

Capra:2001:MMC

[CEM01] L. Capra, W. Emmerich, and C. Mascolo. Middleware for mobile computing: Awareness vs. transparency. In IEEE [IEE01], page 164. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Candea:2001:RRT

[CF01] G. Candea and A. Fox. Recursive restartability: Turning the reboot sledgehammer into a scalpel. In IEEE [IEE01], pages 125–132. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Carter:1993:FTB

[CFH^{+}93] John B. Carter, Bryan Ford, Mike Hibler, Ravindra Kuramkote, Jeffrey Law, Jay Lepreau, Douglas B. Orr, Leigh Stoller, and Mark Swanson. FLEX: a tool for building efficient and flexible systems. In IEEE [IEE93], pages 198–202. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL

ftp://mancos.cs.utah.edu//
papers/flex-wwos4.ps.Z;
http://ieeexplore.ieee.org/
iel2/918/8054/00348148.pdf.

IEEE catalog number 93TH0553-8.

Chase:1993:SIS

[CFL93] Jeffrey S. Chase, Michael J. Feeley, and Henry M. Levy. Some issues for single address space systems. In IEEE [IEE93], pages 150–154. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348157.pdf. IEEE

catalog number 93TH0553-8.

Chapin:1997:FLM

[Cha97]  J. Chapin.  A fresh look

at memory hierarchy management. In IEEE [IEE97], pages 130–134. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Chen:1993:SMS

[Che93] J. Bradley Chen. Software methods for system address tracing. In IEEE [IEE93], pages 178–185. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348152.pdf. IEEE

catalog number 93TH0553-8.

Chiueh:1997:DSS

[Chi97] T. Chiueh. Distributed systems support for networked games. In IEEE [IEE97], pages 99–105. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Cabrera:2001:HAG

[CJT01] L. F. Cabrera, M. B. Jones, and M. Theimer. Herald: Achieving a global event notification service. In IEEE [IEE01], pages 87–94. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Coady:2001:AOS

[CKF^{+}01] Y. Coady, G. Kiczales, M. Feeley, N. Hutchinson, J. S. Ong, and S. Gudmundson. Aspectoriented system structure. In IEEE [IEE01], page 166. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Carter:1995:DSM

[CKK95] J. B. Carter, D. Khandekar, and L. Kamb. Distributed shared memory: Where we are and where we should be headed. In IEEE [IEE95], pages 119–123. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Chen:1999:RHa

[CL99a] P. Chen and D. Lowell. Reliability hierarchies. In IEEE [IEE99], pages 168–173. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Chen:1999:RHb

[CL99b] P. Chen and D. Lowell. Reliability hierarchies. In IEEE [IEE99], pages 168–173. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Chase:1992:OSA

[CLBHL92] J. Chase, H. Levy, M. Baker-Harvey, and E. Lazowska. Opal: a single address space system for 64-bit architecture address space. In IEEE [IEE92], pages 80–85. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275684.pdf. IEEE

catalog number 92TH0413-5.

Chen:2001:WVB

[CN01] P. M. Chen and B. D. Noble. When virtual is better than real. In IEEE [IEE01], pages 133–138. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Collins:1992:SIM

[Col92]  K. R. Collins.  Support-

ing the information mesh. In IEEE [IEE92], pages 56– 61. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275688.pdf. IEEE

catalog number 92TH0413-5.

Castro:2001:UAI

[CRL01] M. Castro, R. Rodrigues, and B. Liskov. Using abstraction to improve fault tolerance. In IEEE [IEE01], pages 27–32. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Campbell:1995:COO

[CT95] R. H. Campbell and S.-M. Tan.

Choices: An object-oriented multimedia operating system. In IEEE [IEE95], pages 90– 94. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Chiueh:1999:IAS

[CVP99] T. Chiueh, G. Venkitachalam, and P. Pradhan. Intra-address space protection using segmentation hardware. In IEEE [IEE99], pages 110–115. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Day:1993:OGM

[Day93]  Mark Day.  Object groups

may be better than pages. In IEEE [IEE93], pages 119– 122. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348161.pdf. IEEE

catalog number 93TH0553-8.

Day:1997:WSG

[Day97] M. Day. What synchronous groupware needs: Notification services. In IEEE [IEE97], pages 118–123. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Draves:1997:OSD

[DBFF97] R. Draves, W. Bolosky, R. Fitzgerald, and C. Fraser. Operating system directions for the next millennium. In IEEE [IEE97], pages 106–111. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Dabek:2001:BPP

[DBK^{+}01] Frank Dabek, Emma Brunskill, M. Frans Kaashoek, David Karger, Robert Morris, Ion Stoica, and Hari Balakrishnan. Building peer-to-peer systems with Chord, a distributed lookup service. In IEEE [IEE01], pages 81–86. ISBN 0-7695-1040-X. US$135.00. URL http:/

                                        /www.pdos.lcs.mit.edu/cgi-
                                        bin/bibtex-entry.cgi?key=

chord:hotos. IEEE catalog number PR01040.

Lara:2001:AAS

[dLWZ01] E. de Lara, D. S. Wallach, and W. Zwaenepoel. Architectures for adaption systems. In IEEE [IEE01], page 167. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Douglis:1992:WWI

[DM92] F. Douglis and B. Marsh.

The workstation as a waystation: integrating mobility into computing environments. In IEEE [IEE92], pages 17– 21. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275695.pdf. IEEE

catalog number 92TH0413-5.

Devarakonda:1995:MSM

[DMK95] M. Devarakonda, A. Mukherjee, and B. Kish. Meta-scripts as a mechanism for complex Web services. In IEEE [IEE95], pages 46–50. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Druschel:1992:MPS

[DPH92] P. Druschel, L. L. Peterson, and N. C. Hutchinson. Modularity and protection should be decoupled. In IEEE [IEE92], pages 95–97. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275681.pdf. IEEE

catalog number 92TH0413-5.

Druschel:1997:EKL

[DPZ97] P. Druschel, V. Pai, and W. Zwaenepoel. Extensible kernels are leading OS research astray. In IEEE [IEE97], pages 38–42. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Druschel:2001:PLS

[DR01] P. Druschel and A. Rowstron. PAST: a large-scale, persistent peer-to-peer storage utility. In IEEE [IEE01], pages 75–80. ISBN 0-7695-1040-X. US$135.00. URL http://

computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Draves:1993:CRT

[Dra93] Richard Draves. The case for run-time replaceable kernel modules. In IEEE [IEE93], pages 160–164. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348155.pdf. IEEE

catalog number 93TH0553-8.

Duchamp:1992:IWM

[Duc92]  D. Duchamp.  Issues in

wireless mobile computing. In IEEE [IEE92], pages 2– 10. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275698.pdf. IEEE

catalog number 92TH0413-5.

Engler:1995:AAL

[EGK95] D. R. Engler, S. K. Gupta, and M. F. Kaashoek. AVM: Application-level virtual memory. In IEEE [IEE95], pages 72–77. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Edwards:2001:SEN

[EH01] A. Edwards and G. Heiser. Secure OS extensibility needn’t cost an arm and a leg. In IEEE [IEE01], page 168. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Engler:1995:EAO

[EK95] D. R. Engler and M. F.

Kaashoek. Exterminate all operating system abstractions. In IEEE [IEE95], pages 78– 85. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Ellis:1999:CHLa

[Ell99a] C. Ellis. The case for higherlevel power management. In IEEE [IEE99], pages 162– 167. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999.

IEEE Computer Society Press order number PR00237.

Ellis:1999:CHLb

[Ell99b] C. Ellis. The case for higherlevel power management. In IEEE [IEE99], pages 162– 167. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Fox:1999:HYSa

[FB99a] A. Fox and E. Brewer. Harvest, yield and scalable tolerant systems. In IEEE [IEE99], pages 174–178. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Fox:1999:HYSb

[FB99b] A. Fox and E. Brewer. Harvest, yield and scalable tolerant systems. In IEEE [IEE99], pages 174–178. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Ford:1997:FTR

[FLCV97] B. Ford, J. Lepreau, S. Clawson, and K. Van Maren. The Flux OS toolkit: Reusable components for OS implementation. In IEEE [IEE97], pages 14–19. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Fleisch:1997:FPG

[Fle97]  B. Fleisch.  The failure

of personalities to generalize. In IEEE [IEE97], pages 8–13. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Flinn:2001:STR

[FNS01] J. Flinn, D. Narayanan, and M. Satyanarayanan. Self-tuned remote execution for pervasive computing. In IEEE [IEE01], pages 61–66. ISBN 0-7695-1040-X. US$135.00. URL

http://www-2.cs.cmu.edu/~
bumba/research/hotos2001-
spectra.pdf;   http://www-
2.cs.cmu.edu/~bumba/research/
hotos2001-spectra.ps.gz. IEEE

catalog number PR01040.

Franz:1997:RTC

[Fra97] M. Franz. Run-time code generation as a central system service. In IEEE [IEE97], pages 112–117. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Faller:1989:PMU

[FS89] Newton Faller and Pedro Salenbauch. PLURIX: a multiprocessing Unix-like operating system. In IEEE [IEE89], pages 29–36. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        267/3322/00109264.pdf. IEEE

catalog number 89TH0281-6.

Forrest:1997:BDC

[FSA97] S. Forrest, A. Somayaji, and D. Ackley. Building diverse computer systems. In IEEE [IEE97], pages 67–72. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Ganger:2001:AC

[Gan01] G. R. Ganger. Authentication confidences. In IEEE [IEE01], page 169. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Grimm:1997:SES

[GB97] R. Grimm and B. Bershad. Security for extensible systems. In IEEE [IEE97], pages 62–66. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Guedes:1993:DSO

[GC93] Paulo Guedes and Miguel Castro. Distributed shared object memory. In IEEE [IEE93], pages 142–149. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348158.pdf. IEEE

catalog number 93TH0553-8.

Grimm:2001:SDP

[GDH^{+}01] R. Grimm, J. Davis, B. Hendrickson, E. Lemar, A. Mac-Beth, S. Swanson, T. Anderson, B. Bershad, G. Borriello, and S. Gribble. Systems directions for pervasive computing. In IEEE [IEE01], pages 147–151. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Golm:2001:BAS

[GKB01] M. Golm, J. Kleinoder, and F. Bellosa. Beyond address spaces-flexibility, performance, protection, and resource management in the type-safe JX operating system. In IEEE [IEE01], pages 3–8. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Guedes:1989:OSS

[GM89] Paulo Guedes and José Alves Marques. Operating system support for an objectoriented environment. In IEEE [IEE89], pages 37–42. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        267/3322/00109265.pdf. IEEE

catalog number 89TH0281-6.

Ganger:2001:BSS

[GN01] G. R. Ganger and D. F. Nagle. Better security via smarter devices. In IEEE [IEE01], pages 100–105. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Gibson:1992:DRD

[GPS92] G. A. Gibson, R. H. Patterson, and M. Satyanarayanan. Disk reads with DRAM latency. In IEEE [IEE92], pages 126–131. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275675.pdf. IEEE

catalog number 92TH0413-5.

Gadde:1997:RRR

[GRC97] S. Gadde, M. Rabinovich, and J. Chase. Reduce, reuse, recycle: An approach to building large Internet caches. In IEEE [IEE97], pages 93– 98. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Gribble:2001:RCS

[Gri01] S. D. Gribble. Robustness in complex systems. In IEEE [IEE01], pages 21–26. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Gwertzman:1995:CGP

[GS95] J. S. Gwertzman and M. Seltzer.

The case for geographical pushcaching. In IEEE [IEE95], pages 51–57. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Goldberg:1993:HPN

[GT93] David Goldberg and Michael M.

Tso. How to program networked portable computers. In IEEE [IEE93], pages 30– 33. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/

                                        918/8054/00348176.pdf. IEEE

catalog number 93TH0553-8.

Hagmann:1989:CWO

[Hag89] Robert Hagmann. Comments on workstation operating systems and virtual memory. In IEEE [IEE89], pages 43–48. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        267/3322/00109266.pdf. IEEE

catalog number 89TH0281-6.

Hagmann:1992:MTV

[Hag92]  R. B. Hagmann.  Medium

term virtual memory replacement. In IEEE [IEE92], pages 142–147. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275672.pdf. IEEE

catalog number 92TH0413-5.

Hamilton:1995:WAR

[Ham95] D. Hamilton. Are we answering the right research questions for commercial operating systems? In IEEE [IEE95], pages 142–144. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Hauser:1992:PIS

[Hau92] C. Hauser. A plea for interfaces that support caching. In IEEE [IEE92], pages 137– 140. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275673.pdf. IEEE

catalog number 92TH0413-5.

Hui:2001:SHS

[HAW^{+}01] K. Hui, J. Appavoo, R. Wisniewski, M. Auslander, D. Edelsohn, B. Gamsa, O. Krieger, B. Rosenburg, and M. Stumm. Supporting hot-swappable components for system software. In IEEE [IEE01], page 170. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Hisgen:1989:ACT

[HBM^{+}89] Andy Hisgen, Andrew Birrell, Timothy Mann, Michael D. Schroeder, and Garret Swart. Availability and consistency tradeoffs in the Echo distributed file system. In IEEE [IEE89], pages 49–54. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109267.pdf. IEEE

catalog number 89TH0281-6.

Herrin:1992:BSR

[HF92] E. H. Herrin II and R. A. Finkel.

The benefits of service rebalancing. In IEEE [IEE92], pages 104–110. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275679.pdf. IEEE

catalog number 92TH0413-5.

Honeyman:1992:LWP

[HHRB92] P. Honeyman, L. Huston, J. Rees, and D. Bachmann. The LITTLE WORK project. In IEEE [IEE92], pages 11– 14. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275697.pdf. IEEE

catalog number 92TH0413-5.

Holland:2001:RIN

[HJM^{+}01] D. A. Holland, W. Josephson, K. Magoutis, M. I. Seltzer, C. A. Stein, and A. Lim. Research issues in no-futz computing. In IEEE [IEE01], pages 106–112. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Hsieh:1993:PRI

[HKW93] Wilson C. Hsieh, M. Frans Kaashoek, and William E. Weihl. The persistent relevance of IPC performance: New techniques for reducing the IPC penalty. In IEEE [IEE93], pages 186–190. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348151.pdf. IEEE

catalog number 93TH0553-8.

Howell:1999:HYG

[HM99] J. Howell and M. Montague.

Hey, you got your compiler in my operating system! In IEEE [IEE99], pages 122– 129. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Howard:1993:URS

[How93] John H. Howard. Using reconciliation to share files between occasionally connected computers. In IEEE [IEE93], pages 56–60. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348172.pdf. IEEE

catalog number 93TH0553-8.

Hutchinson:1989:XKO

[HPR89] Norman C. Hutchinson, Larry L.

Peterson, and Herman Rao. The x-kernel: an open operating system design. In IEEE [IEE89], pages 55–59. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109268.pdf. IEEE

catalog number 89TH0281-6.

IEEE:1987:PWW

[IEE87]  IEEE, editor.  Proceedings

/ Workshop on Workstation Operating Systems, November 5–6, 1987, Cambridge, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1987. ISBN ???? LCCN ????

IEEE:1989:WOS

[IEE89]  IEEE, editor.  Workstation

operating systems: proceedings of the Second Workshop on Workstation Operating Systems (WWOS-II), September 27–29, 1989, Pacific Grove, CA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1989. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. IEEE catalog number 89TH0281-6.

IEEE:1992:TWW

[IEE92] IEEE, editor. Third Workshop on Workstation Operating Systems, April 23–24, 1992, Key Biscayne, Florida: proceedings. IEEE Computer Society Press, Los Alamitos, Calif., 1992. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. IEEE catalog number 92TH0413-5.

IEEE:1993:FIW

[IEE93] IEEE, editor. Fourth IEEE Workshop on Workstation Operating Systems (WWOS-IV), October 14–15, 1993, Napa, CA. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1993. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. IEEE catalog number 93TH0553-8.

IEEE:1995:FWH

[IEE95] IEEE, editor. Fifth Workshop on Hot Topics in Operating Systems (HotOS-V): proceedings, May 4– 5, 1995, Orcas Island, Washington. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1995. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

IEEE:1997:SWH

[IEE97] IEEE, editor. The Sixth Workshop on Hot Topics in Operating Systems: May 5–6, 1997, Cape Cod, Massachusetts. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1997. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

IEEE:1999:SWH

[IEE99] IEEE, editor. The Seventh Workshop on Hot Topics in Operating Systems: [HotOS-VII]: 29–30 March 1999, Rio Rico, Arizona. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 1999. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

IEEE:2001:EIW

[IEE01] IEEE, editor. Eighth IEEE Workshop on Hot Topics in Operating Systems (HotOS-VIII). May 20–23, 2001, Schloss Elmau, Germany. IEEE Computer Society Press, 1109 Spring Street, Suite 300, Silver Spring, MD 20910, USA, 2001. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Itoi:1999:PSSa

[IH99a] N. Itoi and P. Honeyman. Practical security systems with Smartcards. In IEEE [IEE99], pages 185–190. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Itoi:1999:PSSb

[IH99b] N. Itoi and P. Honeyman. Practical security systems with Smartcards. In IEEE [IEE99], pages 185–190. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Ingram:1999:SRT

[Ing99] D. Ingram. Soft real time scheduling for general purpose client-server systems. In IEEE [IEE99], pages 130– 135. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Jeffay:1992:KSR

[Jef92] K. Jeffay. On kernel support for real-time multimedia applications. In IEEE [IEE92], pages 39–46. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275691.pdf. IEEE

catalog number 92TH0413-5.

Jaeger:1999:FACa

[JELP99a] T. Jaeger, K. Elphinstone, J. Liedtke, and V. Panteleenko. Flexible access control using IPC redirection. In IEEE [IEE99], pages 191–196. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Jaeger:1999:FACb

[JELP99b] T. Jaeger, K. Elphinstone, J. Liedtke, and V. Panteleenko. Flexible access control using IPC redirection. In IEEE [IEE99], pages 191–196. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Juul:1995:MAC

[JF95] N. C. Juul and B. D. Fleisch. A memory approach to consistent, reliable distributed shared memory. In IEEE [IEE95], pages 108–112. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Jones:1995:MRT

[JLDB95] M. B. Jones, P. J. Leach, R. P. Draves, and J. S. Barrera. Modular real-time resource management in the Rialto operating system. In IEEE [IEE95], pages 12–17. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Johnson:1993:UMH

[Joh93] David B. Johnson. Ubiquitous mobile host internetworking. In IEEE [IEE93], pages 85–90. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348167.pdf. IEEE

catalog number 93TH0553-8.

Jones:1992:TIU

[Jon92] M. B. Jones. Transparently interposing user code at the system interface. In IEEE [IEE92], pages 98–103. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275680.pdf. IEEE

catalog number 92TH0413-5.

Jones:1993:ART

[Jon93] Michael B. Jones. Adaptive realtime resource management supporting composition of independently authored time-critical services. In IEEE [IEE93], pages 135–139. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348159.pdf. IEEE

catalog number 93TH0553-8.

Jones:1999:PYH

[JR99] M. Jones and J. Regehr. The problems you’re having may not be the problems you think you’re having: Results from a latency study of Windows NT. In IEEE [IEE99], pages 96–101. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Johansen:1995:OSS

[JVS95] D. Johansen, R. Van Renesse, and F. B. Schneider. Operating system support for mobile agents. In IEEE [IEE95], pages 42–45. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Keeton:2001:BAC

[KA01] K. Keeton and E. Anderson. A backup appliance composed of high-capacity disk drives. In IEEE [IEE01], page 171. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Kazar:1989:URS

[Kaz89] Michael Leon Kazar. Ubik: replicated servers made easy. In IEEE [IEE89], pages 60– 67. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL

http://ieeexplore.ieee.org/
iel5/267/3322/00109269.pdf.

IEEE catalog number 89TH0281-6.

Kidston:1999:TCM

[KBK99] D. Kidston, J. Black, and T. Kunz. Transparent communication management in wireless networks. In IEEE [IEE99], pages 64–69. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Katseff:1992:CIW

[KGLR92] H. P. Katseff, R. D. Gaglianello, T. B. London, and B. S. Robinson. Communications-intensive workstations. In IEEE [IEE92], pages 22–28. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275694.pdf. IEEE

catalog number 92TH0413-5.

Katcher:1995:RTO

[KKS95] D. I. Katcher, K. A. Kettler, and J. K. Strosnider. Real-time operating systems for multimedia processing. In IEEE [IEE95], pages 18–23. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Kroeger:1999:CEF

[KL99] T. Kroeger and D. Long. The case for efficient file access pattern modeling. In IEEE [IEE99], pages 14–19. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Kiczales:1993:NCO

[KLM^{+}93] Gregor Kiczales, John Lamping, Chris Maeda, David Keppel, and Dylan McNamee. The need for customizable operating systems. In IEEE [IEE93], pages 165–169. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348154.pdf. IEEE

catalog number 93TH0553-8.

Kiciman:2001:TZC

[KMF01] E. Kiciman, L. Melloul, and A. Fox. Towards zero-code service composition. In IEEE [IEE01], page 172. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Kumar:1993:SAS

[KS93] Puneet Kumar and M. Satyanarayanan. Supporting applicationspecific resolution in an optimistically replicated file system. In IEEE [IEE93], pages 66–70. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348170.pdf. IEEE

catalog number 93TH0553-8.

Khalidi:1993:VMS

[KTNW93] Yousef A. Khalidi, Madhusudhan Talluri, Michael N. Nelson, and Dock Williams. Virtual memory support for multiple page sizes. In IEEE [IEE93], pages 104–109. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL

http://ieeexplore.ieee.org/

iel2/918/8054/00348164.pdf.

IEEE catalog number 93TH0553-8.

Liedtke:1997:AIP

[LESH97] J. Liedtke, K. Elphinstone, S. Schoenberg, and H. Haertig. Achieved IPC performance. In IEEE [IEE97], pages 28–31. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Loser:2001:SIR

[LHR01] J. Loser, H. Hartig, and L. Reuther. A streaming interface for real-time interprocess communication. In IEEE [IEE01], page 174. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Liedtke:1997:PDS

[LIJ97] J. Liedtke, N. Islam, and T. Jaeger. Preventing denialof-service attacks on a ?-kernel for WebOSes. In IEEE [IEE97], pages 73–81. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Lauvset:2001:SMM

[LJJM01] K. J. Lauvset, K. Jacobsen, D. Johansen, and K. Marzullo. Separating mobility from mobile agents. In IEEE [IEE01], page 173. ISBN 0-7695-1040-X. US$135.00. URL http://

                                        computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Landay:1993:UII

[LK93] James A. Landay and Todd R. Kaufmann. User interface issues in mobile computing. In IEEE [IEE93], pages 40– 47. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348174.pdf. IEEE

catalog number 93TH0553-8.

Livny:1992:MWF

[LL92] M. Livny and M. Litzkow. Making workstations a friendly environment for batch jobs. In IEEE [IEE92], pages 62– 67. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275687.pdf. IEEE

catalog number 92TH0413-5.

Leung:1989:CAD

[LMMW89] W. H. Leung, L. F. Morgan, M. J. Morgan, and B. F. Wong. The connector and active devices mechanisms for constructing multimedia applications. In IEEE [IEE89], pages 68–72. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109270.pdf. IEEE

catalog number 89TH0281-6.

Loucks:1993:MBO

[LMR93] Larry K. Loucks, Ravi Manikundalam, and Freeman L. Rawson III. A microkernel-based operating system for personal digital assistants. In IEEE [IEE93], pages 9–13. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348180.pdf. IEEE

catalog number 93TH0553-8.

Lara:1999:CDA

[LPTL99] E. De Lara, K. Petersen, D. Terry, and A. LaMarca. Caching documents with active properties. In IEEE [IEE99], pages 8–13. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Lindstroem:1995:GUT

[LRD95] A. Lindstroem, J. Rosenberg, and A. Dearle. The grand unified theory of address spaces. In IEEE [IEE95], pages 66– 71. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Liskov:1992:EDB

[LS92] B. Liskov and L. Shrira. Escaping the disk bottleneck in fast transaction processing. In IEEE [IEE92], pages 118–121. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275677.pdf. IEEE

catalog number 92TH0413-5.

Lu:1995:IDC

[LS95] Q. Lu and M. Satyanarayanan. Improving data consistency in mobile computing using isolation-only transactions. In IEEE [IEE95], pages 124– 128. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Liskov:1989:EMO

[LSW89] Barbara Liskov, Liuba Shrira, and John Wroclawski. Efficient at-most-once messages based on synchronized clocks. In IEEE [IEE89], pages 73–81. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        267/3322/00109271.pdf. IEEE

catalog number 89TH0281-6.

Liedtke:1999:HSU

[LUEJ99] J. Liedtke, V. Uhlig, K. Elphinstone, and T. Jaeger. How to schedule unlimited memory pinning of untrusted processes or provisional ideas about serviceneutrality. In IEEE [IEE99], pages 153–161. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Leach:1997:QRA

[LW97] P. Leach and C. Weider. Query routing: Applying systems thinking to Internet search. In IEEE [IEE97], pages 82– 86. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Liedtke:2001:LPS

[LW01] J. Liedtke and H. Wenske. Lazy process switching. In IEEE [IEE01], pages 15–20. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Lea:1995:AOS

[LYI95] R. Lea, Y. Yokote, and J.- I. Itoh. Adaptive operating system design using reflection. In IEEE [IEE95], pages 95–100. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Maeda:1993:SS

[MB93] Chris Maeda and Brian N. Bershad. Service without servers. In IEEE [IEE93], pages 170–176. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348153.pdf. IEEE

catalog number 93TH0553-8.

Miller:2001:HHP

[MBL01] E. L. Miller, S. A. Brandt, and D. D. E. Long. HeRMES: Highperformance reliable MRAMenabled storage. In IEEE [IEE01], pages 95–99. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Mendelsohn:1997:OSC

[Men97] N. Mendelsohn. Operating systems for component software environments. In IEEE [IEE97], pages 49–55. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

McJones:1987:TSD

[MH87] P. McJones and A. Hisgen.

The Topaz system: Distributed multiprocessor personal computing. In IEEE [IEE87], page ?? ISBN ???? LCCN

???? URL http://citeseer.
nj.nec.com/mcjones87topaz.
html.

Mazieres:1997:SAN

[MK97] D. Mazieres and M. Kaashoek.

Secure applications need flexible operating systems. In IEEE [IEE97], pages 56–61. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Montz:1995:SCO

[MMOP95] A. B. Montz, D. Mosberger, S. W. O’Malley, and L. L. Peterson. Scout: a communicationsoriented operating system. In IEEE [IEE95], pages 58–61. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Michiels:2001:DUA

[MMWV01] S. Michiels, F. Matthijs, D. Walravens, and P. Verbaeten. DiPS: a unifying approach for developing system software. In IEEE [IEE01], page 175. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Mogul:1989:WRA

[Mog89] Jeffrey C. Mogul. What is the right amount of statelessness in a file server? In IEEE [IEE89], pages 82–85. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        267/3322/00109272.pdf. IEEE

catalog number 89TH0281-6.

Mogul:1992:SLU

[Mog92]  J. C. Mogul.  SPECmarks

are leading us astray. In IEEE [IEE92], pages 160– 161. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275669.pdf. IEEE

catalog number 92TH0413-5.

Mogul:1993:BMD

[Mog93] Jeffrey C. Mogul. Big memories on the desktop. In IEEE [IEE93], pages 110– 115. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348163.pdf. IEEE

catalog number 93TH0553-8.

Mogul:1999:BMO

[Mog99]  J. Mogul.  Brittle metrics

in operating systems research.

In IEEE [IEE99], pages 90– 95. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Mogul:2001:TRD

[Mog01] J. C. Mogul. Toward a rigorous data type model for HTTP. In IEEE [IEE01], page 176. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Mosberger:1993:CPH

[MP93] David Mosberger and Larry L.

Peterson. Careful protocols or how to use highly reliable networks. In IEEE [IEE93], pages 80–84. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348168.pdf. IEEE

catalog number 93TH0553-8.

Mazieres:2001:DTY

[MS01] D. Mazieres and D. Shasha.

Don’t trust your file server. In IEEE [IEE01], pages 113– 118. ISBN 0-7695-1040-X. US$135.00. URL http://

computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Mah:1993:PNV

[MSK^{+}93] Bruce A. Mah, Srinivasan Seshan, Kimberly Keeton, Randy H. Katz, and Domenico Ferrari.

Providing network video service to mobile clients. In IEEE [IEE93], pages 48–54. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL

http://ieeexplore.ieee.org/
iel2/918/8054/00348173.pdf.

IEEE catalog number 93TH0553-8.

Mercer:1993:PCR

[MST93] Clifford W. Mercer, Stefan Savage, and Hideyuki Tokuda. Processor capacity reserves: An abstraction for managing processor usage. In IEEE [IEE93], pages 129–134. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348160.pdf. IEEE

catalog number 93TH0553-8.

Mullender:1987:AWO

[Mul87] S. J. Mullender. Amoeba as a workstation operating system. In IEEE [IEE87], page ?? ISBN ???? LCCN ????

Myers:1993:RIP

[Mye93] Andrew C. Myers. Resolving the integrity/performance conflict. In IEEE [IEE93], pages 156–159. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348156.pdf. IEEE

catalog number 93TH0553-8.

Neufeld:1989:RPD

[NC89] Gerald Neufeld and Samuel Chanson. The Raven project. In IEEE [IEE89], pages 86– 90. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL

http://ieeexplore.ieee.org/
iel5/267/3322/00109273.pdf.

IEEE catalog number 89TH0281-6.

Neuman:1989:WVS

[Neu89] B. Clifford Neuman. Workstations and the virtual system model. In IEEE [IEE89], pages 91–95. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109274.pdf. IEEE

catalog number 89TH0281-6.

Nilsen:1995:RTN

[Nil95] K. Nilsen. Real-time is no longer a small specialized niche. In IEEE [IEE95], pages 2– 7. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Nelson:1993:GSC

[NK93] Michael N. Nelson and Yousef A.

Khalidi. Generic support for caching and disconnected operation. In IEEE [IEE93], pages 61–65. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348171.pdf. IEEE

catalog number 93TH0553-8.

Noble:1999:CBT

[NLP99] B. Noble, L. Li, and A. Prakash. The case for better throughput estimation. In IEEE [IEE99], pages 70–75. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Neugebauer:2001:EJA

[NM01] R. Neugebauer and D. McAuley. Energy is just another resource: Energy accounting and energy pricing in the Nemesis OS. In IEEE [IEE01], pages 67–74. ISBN 0-7695-1040-X. US$135.00. URL http://

                                        computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Narten:1992:RMR

[NY92] T. Narten and R. Yavagkar. Remote memory as a resource in distributed systems. In IEEE [IEE92], pages 132– 136. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275674.pdf. IEEE

catalog number 92TH0413-5.

Ordille:1992:LLW

[OM92] J. J. Ordille and B. P. Miller. Lost in a labyrinth of workstations. In IEEE [IEE92], pages 52–55. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275689.pdf. IEEE

catalog number 92TH0413-5.

Orr:1995:AMP

[Orr95] D. B. Orr. Application of metaprotocols to improve OS services. In IEEE [IEE95], pages 101–107. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

OQuin:1989:ATO

[OS89] J. C. O’Quin and C. H.

Sauer. AIX 3 technology (operating system). In IEEE [IEE89], pages 96–100. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109275.pdf. IEEE

catalog number 89TH0281-6.

OToole:1995:SDM

[OS95] J. O’Toole and L. Shrira.

Shared data management needs adaptive methods. In IEEE [IEE95], pages 129–135. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Pasquale:1992:SDI

[Pas92] J. Pasquale. I/O system design for intensive multimedia I/O. In IEEE [IEE92], pages 29–33. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        891/6831/00275693.pdf. IEEE

catalog number 92TH0413-5.

Pradhan:1999:OSS

[PC99] P. Pradhan and T. Chiueh. Operating systems support for programmable cluster-based interact routers. In IEEE [IEE99], pages 76–81. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Perez-Cortes:1995:GSS

[PCDH95] E. Perez-Cortes, P. Dechamboux, and J. Han. Generic support for synchronization and consistency in Arias. In IEEE [IEE95], pages 113–118. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Pu:1993:FBS

[PF93] Calton Pu and Robert M. Fuhrer. Feedback-based scheduling: a toolbox approach. In IEEE [IEE93], pages 124– 128. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348177.pdf. IEEE

catalog number 93TH0553-8.

Pradhan:2001:DIS

[PGC01] P. Pradhan, K. Gopalan, and T.-C. Chiueh. Design issues in system support for programmable routers. In IEEE [IEE01], pages 9–14. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Peterson:1999:SGP

[PKL99] L. Peterson, S. Karlin, and K. Li.

OS support for general-purpose routers. In IEEE [IEE99], pages 38–43. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Preguica:2001:SDO

[PMDD01] N. Preguica, J. Martins, H. Domingos, and S. Duarte. Supporting disconnected operation in DOORS. In IEEE [IEE01], page 179. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Petrou:2001:HGS

[PNG^{+}01] D. Petrou, D. Narayanan, G. R. Ganger, G. A. Gibson, and E. Shriver. Hinting for goodness’ sake. In IEEE [IEE01], page 177. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Piumarta:2001:AVK

[POBF01] I. Piumarta, F. Ogel, C. Baillarguet, and B. Folliot. Applying the VVM kernel to flexible Web caches. In IEEE [IEE01], page 178. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Pucci:1992:DCA

[Puc92] M. F. Pucci. Device control abstractions for workstation operating system. In IEEE [IEE92], pages 34–38. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275692.pdf. IEEE

catalog number 92TH0413-5.

Rawson:1997:EDM

[Raw97] F. Rawson. Experience with the development of a microkernelbased, multi-server operating system. In IEEE [IEE97], pages 2–7. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Rashid:1989:MFO

[RBF^{+}89] Richard Rashid, Robert Baron, Alessandro Forin, David Golub, Michael Jones, Daniel Orr, and Richard Sanzi. Mach: a foundation for open systems (operating systems). In IEEE [IEE89], pages 109–113. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109278.pdf. IEEE

catalog number 89TH0281-6.

Razafimahefa:2001:LAW

[RBP01] C. Razafimahefa, C. Bryce, and M. Pawlak. The Lana approach to wireless computing. In IEEE [IEE01], page 180. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Rahul:1999:ESA

[RBS99] H. Rahul, H. Balakrishnan, and S. Seshan. An end-system architecture for unified congestion management. In IEEE [IEE99], pages 52–57. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Redell:1992:WAD

[Red92]  D. D. Redell.  Worksta-

tion autonomy is a dead issue. In IEEE [IEE92], pages 15–16. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275696.pdf. IEEE

catalog number 92TH0413-5.

Ramakrishnan:1989:WNP

[RH89] K. K. Ramakrishnan and William R. Hawe. The workstation on the network: performance considerations for the communications interface. In IEEE [IEE89], pages 105–108. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL

http://ieeexplore.ieee.org/
iel5/267/3322/00109277.pdf.

IEEE catalog number 89TH0281-6.

Ricci:2001:APA

[RL01] R. Ricci and J. Lepreau. Active protocols for agile censorresistant networks. In IEEE [IEE01], pages 119–124. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Rowstron:2001:PMR

[RLB01] A. I. T. Rowstron, N. Lawrence, and C. M. Bishop. Probabilistic modelling of replica divergence. In IEEE [IEE01], pages 55–60. ISBN 0-7695-1040-X. US$135.00. URL http://

computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Reed:1999:XAE

[RPEM99] D. Reed, I. Pratt, S. Early, and P. Menage. Xenoservers: Accountable execution of untrusted programs. In IEEE [IEE99], pages 136–141. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Schilit:1993:PMC

[SAG^{+}93] Bill N. Schilit, Norman Adams, Rich Gold, Michael M. Tso, and Roy Want. The ParcTab mobile computing system. In IEEE [IEE93], pages 34–39. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL

http://ieeexplore.ieee.org/
iel2/918/8054/00348175.pdf.

IEEE catalog number 93TH0553-8.

Satyanarayanan:1989:CHA

[Sat89]  M. Satyanarayanan.  Coda:

a highly available file system for a distributed workstation environment. In IEEE [IEE89], pages 114–116. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109279.pdf. IEEE

catalog number 89TH0281-6.

Snoeren:2001:RIM

[SBK01] A. C. Snoeren, H. Balakrishnan, and M. F. Kaashoek. Reconsidering internet mobility. In IEEE [IEE01], pages 41–46. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Savage:1999:CIT

[SCA99] S. Savage, N. Cardwell, and T. Anderson. The case for informed transport protocols. In IEEE [IEE99], pages 58– 63. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Schatz:1989:PIE

[Sch89] Bruce R. Schatz. A prototype information environment. In IEEE [IEE89], pages 118–124. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        267/3322/00109280.pdf. IEEE

catalog number 89TH0281-6.

Santry:1999:EFS

[SFHV99] D. Santry, M. Feeley, N. Hutchinson, and A. Veitch. Elephant: The file system that never forgets. In IEEE [IEE99], pages 2–7. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Scott:1992:SMO

[SG92] M. L. Scott and W. Garnett. Shared memory ought to be commonplace. In IEEE [IEE92], pages 86–90. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275683.pdf. IEEE

catalog number 92TH0413-5.

Sullivan:1999:TCR

[SHS99] D. Sullivan, R. Haas, and M. Seltzer. Tickets and currencies revisited: Extensions to multi-resouce lottery scheduling. In IEEE [IEE99], pages 148–152. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Satyanarayanan:1992:ULD

[SKKM92] M. Satyanarayanan, J. J. Kistler, P. Kumar, and H. Mashburn. On the ubiquity of logging in distributed file systems. In IEEE [IEE92], pages 122–125. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275676.pdf. IEEE

catalog number 92TH0413-5.

Satyanarayanan:1987:CRD

[SKS87] M. Satyanarayanan, J. J. Kistler, and E. H. Siegel. Coda: a resilient distributed file system. In IEEE [IEE87], page

????  ISBN ????  LCCN

????  URL http://www.

                                        cs.cmu.edu/afs/cs/project/
                                        coda/Web/docs-coda.html.

Seltzer:1999:CAS

[SKSZ99] M. Seltzer, D. Krinsky, K. Smith, and X. Zhang. The case for application-specific benchmarking. In IEEE [IEE99], pages 102–109. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Scott:1989:MUM

[SLM89] Michael L. Scott, Thomas J. LeBlanc, and Brian D. Marsh. A multi-user, multi-language open operating system. In IEEE [IEE89], pages 125–129. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

                                        /ieeexplore.ieee.org/iel5/
                                        267/3322/00109281.pdf. IEEE

catalog number 89TH0281-6.

Stratford:1999:EAA

[SM99] N. Stratford and R. Mortier. An economic approach to adaptive resource management. In IEEE [IEE99], pages 142– 147. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Stanley-Marbell:2001:SMS

[SMBNI01] P. Stanley-Marbell, C. Borcea, K. Nagaraja, and L. Iftode.

Smart messages: a system architecture for large networks of embedded systems. In IEEE [IEE01], page 181. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Satyanarayanan:1997:RTM

[SN97] M. Satyanarayanan and B. Noble. The role of trace modulation in building mobile computing systems. In IEEE [IEE97], pages 135–139. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Steere:1995:UDS

[SS95] D. Steere and M. Satyanarayanan. Using dynamic sets to overcome high I/O latencies during search. In IEEE [IEE95], pages 136–141. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Seltzer:1997:SMS

[SS97] M. Seltzer and C. Small. Selfmonitoring and self-adapting operating systems. In IEEE [IEE97], pages 124–129. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

Sultan:2001:TLS

[SSI01] F. Sultan, K. Srinivasan, and L. Iftode. Transport layer support for highly-available network services. In IEEE [IEE01], page 182. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Stankovic:1995:CMS

[Sta95]  J. A. Stankovic.  Continu-

ous and multimedia OS support in real-time control applications. In IEEE [IEE95], pages 8–11. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Theimer:1993:OSI

[TDW93] Marvin Theimer, Alan J. Demers, and Brent B. Welch. Operating system issues for PDAs. In IEEE [IEE93], pages 2–8. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348181.pdf. IEEE

catalog number 93TH0553-8.

Tol:1999:FSC

[TI99] R. Tol and P. Ijdens. File systems for CE equipment.

In IEEE [IEE99], pages 20– 23. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Tsai:1999:CMSa

[TRP99a] M. Tsai, P. Reiher, and G. Popek. Command management system for next-generation user input. In IEEE [IEE99], pages 179–184. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Tsai:1999:CMSb

[TRP99b] M. Tsai, P. Reiher, and G. Popek. Command management system for next-generation user input. In IEEE [IEE99], pages 179–184. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Tullmann:1997:FMP

[TTML97] P. Tullmann, J. Turner, J. Mc-Corquodale, and J. Lepreau. Formal methods: a practical tool for OS implementors. In IEEE [IEE97], pages 20–27. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.

VanDoorn:1995:PEO

[VHT95] L. Van Doorn, P. Homburg, and A. S. Tanenbaum. Paramecium: An extensible objectbased kernel. In IEEE [IEE95], pages 86–89. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Noordende:2001:MDM

[vNBTvS01] G. v. Noordende, F. Brazier, A. S. Tanenbaum, and M. van Steen. Mansion, A distributed multi-agent system. In IEEE [IEE01], page 183. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

vanRenesse:1987:WCR

[vRTS87] R. van Renesse, A. S. Tanenbaum, and G. J. Sharp. The workstation: Computing resource or just a terminal? In IEEE [IEE87], page ?? ISBN ???? LCCN ????

Veitch:2001:TGS

[VRTW01] A. Veitch, E. Riedel, S. Towers, and J. Wilkes. Towards global storage management and data placement. In IEEE [IEE01], page 184. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Wang:1993:XWA

[WA93] Randolph Wang and Thomas E.

Anderson. xFS: a wide area mass storage file system. In IEEE [IEE93], pages 71–78. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

/ieeexplore.ieee.org/iel2/
918/8054/00348169.pdf. IEEE

catalog number 93TH0553-8.

Watson:1995:EWC

[Wat95] T. Watson. Effective wireless communication through application partitioning. In IEEE [IEE95], pages 24–27. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Welch:1989:SPS

[WBD^{+}89] Brent Welch, Mary Baker, Fred Douglis, John Hartman, Mendel Rosenblum, and John Ousterhout. Sprite position statement: use distributed state for failure recovery. In IEEE [IEE89], pages 130–133. ISBN 0-8186-2003-X, 0-8186-5003-6 (microfiche). LCCN QA76.76.O63 W667 1989. URL http:/

/ieeexplore.ieee.org/iel5/
267/3322/00109282.pdf. IEEE

catalog number 89TH0281-6.

Welsh:2001:VCH

[WC01] M. Welsh and D. Culler. Virtualization considered harmful:

OS design directions for wellconditioned services. In IEEE [IEE01], pages 139–146. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Welch:1992:PPD

[Wel92]  B. Welch.  A proposal for

a public domain systems library. In IEEE [IEE92], pages 111–115. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275678.pdf. IEEE

catalog number 92TH0413-5.

Wilkes:2001:EDR

[WGJ^{+}01] J. Wilkes, P. Goldsack, G. Janakiraman, L. Russell, S. Singhal, and A. Thomas. eOS — the dawn of the resource economy. In IEEE [IEE01], page 188. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Warfield:2001:IGP

[WH01] A. Warfield and N. Hutchinson. The importance of good plumbing-reconsidering infrastructure in distributed systems. In IEEE [IEE01], page 187. ISBN 0-7695-1040-X. US$135.00. URL http://

computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Wilkes:1992:CNA

[Wil92] M. V. Wilkes. The case for a new approach to operating systems for personal computers and work stations. In IEEE [IEE92], pages 164–167. ISBN 0-8186-2555-4 (paper), 0-8186-2556-2 (microfiche). LCCN QA 76.76 O63 W667 1992. URL http:/

/ieeexplore.ieee.org/iel2/
891/6831/00275668.pdf. IEEE

catalog number 92TH0413-5.

Waldman:2001:CRP

[WM01] M. Waldman and D. Mazieres.

Censorship resistant publishing through document entanglements. In IEEE [IEE01], page 185. ISBN 0-7695-1040-X. US$135.00. URL http://

computer.org/CSPRESS/CATALOG/

pr01040.htm. IEEE catalog number PR01040.

Wang:2001:CFS

[WRPK01] A.-I. A. Wang, P. Reiher, G. J. Popek, and G. H. Kuenning. The Conquest file system — life after disks. In IEEE [IEE01], page 186. ISBN 0-7695-1040-X. US$135.00.

URL http://computer.org/
CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Wu:1993:ENM

[WZ93] Michael Wu and Willy Zwaenepoel.

eNVy: a nonvolatile main memory storage system. In IEEE [IEE93], pages 116– 118. ISBN 0-8186-4000-6 (paper), 0-8186-4001-4 (microfiche). LCCN QA76.76.O63 W667 1993. URL http:/

                                        /ieeexplore.ieee.org/iel2/
                                        918/8054/00348162.pdf. IEEE

catalog number 93TH0553-8.

Yocum:2001:ACP

[YCV01] K. Yocum, J. Chase, and A. Vahdat. Anypoint communication protocol. In IEEE [IEE01], pages 189–190. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Young:2001:PIP

[YLS^{+}01] Cliff Young, Y. N. Lakshman, Tom Szymanski, John Reppy, David Presotto, Rob Pike, Girija Narlikar, Sape Mullender, and Eric Grosse. Protium, an infrastructure for partitioned applications. In IEEE [IEE01], pages 47–54. ISBN 0-7695-1040-X. US$135.00.

                                        URL http://computer.org/
                                        CSPRESS/CATALOG/pr01040.htm.

IEEE catalog number PR01040.

Yarvis:1999:CFD

[YRP99] M. Yarvis, P. Reiher, and G. Popek. Conductor: a framework for distributed adaptation. In IEEE [IEE99], pages 44–51. ISBN 0-7695-0237-7, 0-7695-0238-5 (case), 0-7695-0239-3 (microfiche). LCCN QA76.76.O63 W6666 1999. IEEE Computer Society Press order number PR00237.

Zenel:1995:ICF

[ZD95] B. Zenel and D. Duchamp. Intelligent communication filtering for limited bandwidth environments. In IEEE [IEE95], pages 28–34. ISBN 0-8186-7081-9 (paper). LCCN QA76.76.O63 W664 1995. IEEE catalog number 95TH8059. IEEE Computer Society Press order number PR07081.

Zenel:1997:GPP

[ZD97] B. Zenel and D. Duchamp. General purpose proxies: Solved and unsolved problems. In IEEE [IEE97], pages 87–92. ISBN 0-8186-7834-8, 0-8186-7835-6 (case), 0-8186-7836-4 (microfiche). LCCN QA76.76.O63 W67 1997. IEEE Computer Society Press order number PR07834. IEEE Order Plan catalog number 97TB8100133.